OpenStudy (anonymous):

How to solve for A,B,C, and D given the following equations: A+C=0 B-D=0 4A+3D+2C=7 A+4B+6C=-6

OpenStudy (agreene):

you would start by solving one of the eqn for any variable, and then plug that information in to a different eqn and solve that for a different var until you eventually have numerical interpretations for the vars.

OpenStudy (phi):

once you have more than 2 variables, I would use elimination. write all equations, with the variables sorted alphabetical in fact, we can leave out the "names" and just write the coefficients

OpenStudy (phi):

A+C=0 1 0 1 0 | 0 B-D=0 0 1 0 -1 | 0 4A+3D+2C=7 4 0 2 3 | 7 A+4B+6C=-6 1 4 6 0 | -6

OpenStudy (anonymous):

oh this is something i can solve with a matrix?

OpenStudy (phi):

yes, now use "elimination" to zero out the first column (if you know how to do that)

OpenStudy (phi):

example: multiply the 1st row by -4 and add to the 3rd row to get a new 3rd row

OpenStudy (phi):

then multiply the 1st row by -1 and add to the 4th row to get a new 4th row

OpenStudy (phi):

you should get 1 0 1 0 | 0 0 1 0 -1 | 0 0 0 -2 3 | 7 0 4 5 0 | -6

OpenStudy (phi):

now zero out the 2nd column below the 2nd row only 1 step: multiply the 2nd row by -4 and add to the 4th row to get a new 4th row

OpenStudy (anonymous):

Well i don't know how to do it on paper but my teacher showed us how to do it on calculator using the matrix and rref so i did it that way. are the answers A=2, B=1, C=-2, and D=1 ?

OpenStudy (phi):

yes, you want to find reduced row, echelon form If you got the correct answer, those values will work in the original equations

OpenStudy (anonymous):

ok. Thanks for the help!

OpenStudy (phi):

yw